### v1

# Dedelayed: Deleting remote inference delay via on-device correction

Abstract
Remote inference allows lightweight devices to leverage powerful cloud models. However, communication network latency makes predictions stale and unsuitable for real-time tasks. To address this, we introduce Dedelayed, a delay-corrective method that mitigates arbitrary remote inference delays, allowing the local device to produce low-latency outputs in real time. Our method employs a lightweight local model that processes the current frame and fuses in features that a heavyweight remote model computes from past frames. On video from the BDD100K driving dataset, Dedelayed improves semantic segmentation accuracy over the stronger of the local-only and remote-only baselines across all realistic communication network delays beyond 33 ms. Without incurring additional delay, it improves accuracy by 6.4 mIoU compared to fully local inference and 9.8 mIoU compared to remote inference, for a round-trip delay of 100 ms. The advantage grows under longer delays and higher-motion scenes, as delay-mitigated split inference sustains accuracy more effectively, providing clear advantages for real-time tasks that must remain aligned with the current world state.
1 Introduction
In soft real-time applications—such as cloud gaming or video conferencing—late outputs may be diminished in value, but are still useful. In these applications, we can offload expensive computations to powerful cloud GPUs to save on-device power. As long as the typical latency is low, the loss of utility from latency is outweighed by power savings and extended battery life. However, in hard real-time applications—such as aerial robotic control or obstacle avoidance—late outputs can be catastrophic, and the system must be designed with a guaranteed deadline. Due to the irreducible high-tail latency in wireless communication, hard real-time applications must be equipped with a fully functional local inference pipeline as a fallback in the case that the remote predictions fail to meet the deadline. In this work, we focus on such real-time applications running on resource-constrained devices, relying on inference using video inputs.
In recent years, various approaches to split computing have been proposed to offload computation of expensive image and video models to the cloud to enable the next-generation of robotic, remote sensing, and wearable technology platforms. For real-time streaming video applications, existing approaches still fall into one or more of three common pitfalls. (1) They allocate all on-device power and computation to a single linear inference pipeline, leaving no resources for a local-only fallback. (2) They do not account for the impact of latency on prediction accuracy. (3) They operate on videos with significantly reduced spatiotemporal resolution to manage computational cost, leaving out rich visual details available from modern camera systems.
Figure 1: Overview of the Dedelayed real-time inference setup. The lightweight local model and a powerful remote model augment each other’s strengths to produce accurate and timely outputs.
To address these limitations, we introduce Dedelayed (Fig.˜1), a co-inference framework that leverages fresh local information to mitigate the effects of remote inference delay. It consists of a local model and a remote model connected over a communication network, and fuses delayed remote signals into a local pipeline operating on fresh on-device inputs, yielding real-time performance that is never worse than either model in isolation. This enables the use of high-capacity cloud models for delay-sensitive applications while avoiding the pitfalls of prior offloading approaches:
  1. 1.
    Full integration with local-only fallback model. No wireless communication channel can offer perfect reliability. For real-time applications with critical deadlines, any remote inference procedure must be accompanied by a lightweight local fallback model. Instead of two redundant inference pipelines, Dedelayed uses a single path based on a local model that optionally incorporates side information from the remote model. We choose a simple method to incorporate this side information—element-wise addition of activation maps—resulting in negligible overhead and well-defined behavior in the absence of remote outputs.
  2. 2.
    Temporal prediction for latency mitigation. During supervised training of the remote model component, we simulate a delay of $D$ frames. In other words, the remote model is trained to predict the future. A delay embedding—similar to a position embedding in text or vision transformers—allows the behavior of the remote model to adapt to changes in the channel. As shown in Fig.˜2, temporally predictive training is able to capture motion dynamics, which can be used to compensate for latency.
  3. 3.
    Mixed-resolution inference. On-device AI video processing at or near the capture resolution and frame rate—typically >1>1 megapixel and >20>20 frames per second—is rarely feasible, even with lightweight models. To save resources, real-time computer vision applications often process each frame independently or use small motion updates instead of natively processing a dense 3D pixel volume. Dedelayed is capable of using a mixed-resolution—instead of reducing the resolution for the entire inference pipeline, only the local model resolution is reduced. In parallel, the remote model operates on multiple, high-resolution frames using a 3D transformer. Thus, the remote model can utilize powerful GPUs to model fine details and motion of delayed frames, while the local component can allocate its resources to modeling the current state of objects and the scene, as shown in Fig.˜3.
Our contributions are threefold:
  1. 1.
    We provide measurements demonstrating how higher degrees of latency hurt the accuracy of dense visual prediction for semantic segmentation of driving scenes.
  2. 2.
    We introduce Dedelayed, a split computing framework for video inputs that integrates the output of a future-predicting remote model with the current output of a local model.
  3. 3.
    Using Dedelayed, we create a video segmentation system for urban driving scenes that outperforms any existing local or remote inference solution, while avoiding the pitfalls that limit the practicality of previous approaches.
For experimental validation, we demonstrate a simple Dedelayed system with an addition-based fusion on off-the-shelf models with minimal architectural changes. This makes it easy to enhance existing pipelines, deploy in practice, and extend to other real-time methods.
Figure 2: To demonstrate the effect of temporally predictive training, we train a 3D transformer to predict the next frame with an MSE loss on pixels. (a) shows the original video frame. (b) shows the difference between (a) and a future frame, with objects such as the traffic sign and road markings in different locations. (c) shows the pixel predictions of the 3D transformer. (d) shows the difference from the true future frame. While the predictive model cannot predict high-frequency details, it is able to accurately model the motion of objects, signs, and road markings.
Figure 3: Example of activation maps from local and remote model components. The remote server uses the higher level of video detail to accurately distinguish and classify objects. The local model provides exact position adjustments based on the current frame. When making predictions from the combined activation map, small details that would be impossible to make out at low resolution (e.g., the distant pedestrians, labeled red) are accurately classified and localized.
2 Background
In the human visual system, the optic nerve can only transmit a small fraction of the information received by the retina (Kelly, 1962). Barlow’s efficient coding hypothesis (Barlow et al., 1961) posits that compression is the primary role of early processing; once this compressed representation is received in deeper layers of the visual cortex, more metabolically intense processing can occur. In the predictive coding model (Rao & Ballard, 1999), this processing is driven by feedback mechanisms that minimize a temporally predictive error signal to create a perceptual model that is consistent with sensory inputs.
Machines equipped with digital video sensors—which are at the center of ongoing innovation in robotics (Kim et al., 2024; O’Neill et al., 2024), remote sensing (Szwarcman et al., 2024; Khani et al., 2021), and wearable technology (Grauman et al., 2022; 2024)—share similar constraints. The throughput and power efficiency of ingesting pixels on the sensor device (e.g., a battery-powered robot) are extremely high—typically tens or hundreds of megapixels per watt-second (Engel et al., 2023). However, moderately sized DNNs can only process visual data at about one megapixel per watt-second (Cai et al., 2023). For more advanced video AI based on autoregressive modeling (Agarwal et al., 2025) or temporal prediction (Assran et al., 2025), the efficiency may be as low as 500 pixels per watt-second. Instead of on-device processing, power constraints can be circumvented by compressing and transmitting video streams to cloud GPU datacenters supported by a 100-megawatt power infrastructure (Goldberg & Kehoe, 2013).
However, fully remote processing is challenging for certain real-time applications (e.g., collision avoidance) due to unreliability in network and cloud infrastructure (Chen et al., 2024; 2025). Thus, delivering predictions by a guaranteed deadline requires a fallback procedure independent of the remote server. In many systems (e.g., autonomous motor vehicles) the limited accuracy and reliability of lightweight local models warrant a human operator as the fallback (Committee, 2021), preventing full automation.
3 Related work
Prior work has extensively explored lightweight architectures (e.g., EfficientViT (Cai et al., 2023), MobileNetV4 (Qin et al., 2024)) that minimize computation to achieve real-time on-device performance. These approaches deliver low latency but are constrained by device power and compute. When devices are too limited, the common alternative is fully remote inference, which offloads computation to servers, but is highly susceptible to network latency.
Split-computing approaches largely focus on distributing workloads rather than optimizing for strict real-time operation. Next-generation compression standards such as MPEG AI (ISO/IEC, 2025) and JPEG AI (Ascenso et al., 2023) target bandwidth reduction via task-specific compression, significantly lowering transmission costs with reasonable compute overhead. However, even they lack explicit mechanisms to anticipate or compensate for network delay, leading to stale predictions misaligned with the current world state.
Other efforts related to our work have also been explored. Clockwork Convnets (Shelhamer et al., 2016) reuse stale features to reduce inference latency, but they offer limited temporal reasoning and operate on a single device. Accel (Jain et al., 2019) warps heavy-model features forward with optical flow and corrects them with a lightweight model, but is also not intended for across-network operation. Adaptive Model Streaming (Khani et al., 2021) streams weight updates from a server to keep a local model fresh, focusing on model adaptation rather than directly mitigating per-frame staleness from communication latency. Though less known, Knowledge Boosting (Srinivas et al., 2024) is very closely related to our work. Like us, it fuses delayed remote features with a small on-device model, but it assumes a fixed delay. We generalize to longer and variable latencies by conditioning on a tunable delay while keeping the design simple and reusable.
4 Method
4.1 Delay mitigation framework
Figure 4: Time progresses left to right. The client-side camera produces video frames, which are sent across a communication network to the server. The server runs a heavyweight model using the latest video frame $x_{t-\tau}$ that it receives, in addition to a context of previously received video frames $x_{<t-\tau}$, as well as the measured delay $\tau$. This produces an output $z_{t-\tau}$ that the server sends to the client. The client pairs the latest received response $z_{t-\tau}$ with a freshly produced video frame $x_{t}$, and runs these inputs through a lightweight model. This finally produces a timely result $\hat{y}_{t}$ that can be used in real-time delay-sensitive applications.
Dedelayed introduces a general framework that improves the accuracy and robustness of real-time inference on resource-constrained sensor devices. It does so by combining the strengths of both local inference and remote inference, while mitigating their weaknesses. The local model has access to the latest sensor data, and yet lacks the computational capability needed to produce accurate outputs. The remote model provides accurate outputs, and yet delivers them with delay. By careful combination of both subsystems, Dedelayed is able to provide bounded performance guarantees—it is never worse than either local inference or remote inference independently. As we will demonstrate later, we are able to glue together the two subsystems in a way that is simple yet effective.
Dedelayed addresses the problem of stale predictions from powerful remote models by integrating them with a lightweight, on-device model. The core idea is to leverage the high-quality features from a heavyweight remote model, despite their inherent delay, by explicitly conditioning them on the measured latency and fusing them early with live information from a local model. This ensures that the final predictions are both accurate and timely.
Dedelayed can be formulated in simple mathematical terms as follows. Given a fresh input frame $x_{t}$ at current time $t$, the final prediction $\hat{y}_{t}$ is computed using a lightweight local model, $f_{\text{light}}$, which processes $x_{t}$ along with time-delayed features $z_{t-\tau}$ from a heavyweight remote model, $f_{\text{heavy}}$. To produce powerful predictive features, the remote model is conditioned on the delay $\tau$, and processes a short clip of past frames $x_{\leq t-\tau}$ ending at time $t-\tau$. This is expressed by the following equations:
$z_{t-\tau}=f_{\text{heavy}}(\tau,\,x_{\leq t-\tau})$ (1)
$\hat{y}_{t}=f_{\text{light}}(x_{t},\,z_{t-\tau})$ (2)
For clarity, the notation is summarized in Table˜1. Fig.˜4 presents a system diagram that demonstrates the fundamental principle we describe, and shows how information propagates through the various subsystems as time progresses.
Table 1: Notation.
Symbol Meaning
$x_{t}$ Input frame at current time $t$
$x_{\leq t-\tau}$ Input frames up to time $t-\tau$
$\hat{y}_{t}$ Prediction for time $t$
$z_{t-\tau}$ Features outputted by heavyweight model run at time $t-\tau$
$\tau$ Delay in time between the old and current frame at time $t$
$D$ Delay in frames between the old and current frame at time $t$
$f_{\text{light}}$ Lightweight model run at time $t$
$f_{\text{heavy}}$ Heavyweight model run at time $t-\tau$
The entire Dedelayed system is trained end-to-end to minimize a task-specific loss function, $\mathcal{L}_{\text{task}}$, evaluated against the ground truth $y_{t}$ for the current frame.
$\mathcal{L}_{\text{task}}=\ell(\hat{y}_{t},y_{t})$
For semantic segmentation, $\ell$ is typically the cross-entropy loss. The objective is to produce predictions $\hat{y}_{t}$ that are accurate at time $t$ on the local device.
In the next section, we detail how we designed a specific implementation to test the Dedelayed framework in action.
4.2 Design and Implementation
Dedelayed consists of a lightweight on-device model and a predictive remote model connected over a communication network. The remote side computes delay-conditioned features from past inputs and returns them to the device, where they are fused early into the local model running on the current input. To provide a guarantee on baseline performance, these components can first be trained to work independently and later fused and trained jointly.
System overview
Information propagates through our system as follows:
  1. 1.
    The local device transmits input frames to the remote via the uplink.
  2. 2.
    Each incoming frame is independently encoded into features using a pretrained 2D ViT backbone on the remote. We maintain a context window of the $K$ most recent features.
  3. 3.
    The $K$ per-frame features are concatenated along the temporal axis, and a learned delay embedding conditioned on the measured delay $\tau$ is added.
  4. 4.
    A 3D ViT encoder followed by learned pooling (MLP–pool–MLP) produces delay-conditioned remote features $z_{t-\tau}$, which are sent back to the device via the downlink.
  5. 5.
    The lightweight local model runs on a fresh input $x_{t}$, and fuses in the remote features $z_{t-\tau}$.
  6. 6.
    The local model finishes decoding the fused representation and outputs labels $\hat{y}_{t}$.
Remote predictive module.
Figure Fig.˜5 visualizes the remote component. The remote model processes a fixed context of $K$ past frames ending at $t-\tau$. Each frame is encoded independently with a 2D ViT—we use EfficientViT-L1 with an effective $8\times 8$ patch. The per-frame features are concatenated along the temporal axis and spatially merged into larger $16\times 16$ patches to keep the sequence length comparable. (When $K=4$, the sequence length is identical.) A learned delay embedding determined by the delay $\tau$ is added, and the result is then processed by a 3D ViT followed by learned pooling (MLP–pool–MLP) to produce delay-conditioned features $z_{t-\tau}$. We pretrain the remote predictive module by attaching a task head to the 3D ViT backbone and training it to predict the target labels $y_{t}$ from inputs up to time $t-\tau$.
Figure 5: Overview of the remote video predictive model, trained to predict the label at index $n$ from a $K$-frame context ending at nDn-D.
Local model and fusion.
The lightweight local model processes the fresh input $x_{t}$. We compute first-stage features $h=\texttt{T1}(x_{t})$ and perform early fusion by element-wise addition with the delayed remote features, $h^{\prime}=h+z_{t-\tau}$. Both tensors have shape $96\times H/8\times W/8$, so no projection or resizing is required. The fused features $h^{\prime}$ are then processed by the remaining local blocks and decoded to the final output $\hat{y}_{t}$. If $z_{t-\tau}$ is unavailable, the local model falls back to $h^{\prime}=h$ and proceeds unchanged.
5 Experiments
We focus our evaluation of Dedelayed on the task of real-time semantic segmentation of driving scenes, a domain where timely and accurate perception is critical for applications such as autonomous driving and robotics. We demonstrate that delay-aware feature fusion can mitigate remote inference latency, sustaining accuracy even when remote predictions are delayed by long communication network latencies.
Experimental setup
Our experiments utilize the BDD100K video dataset (Yu et al., 2020), containing video of driving scenes at 30 frames per second (fps). Since the dataset does not provide dense segmentation labels for all video frames, we generated pseudo-labels using a pretrained EoMT (Kerssies et al., 2025) model, ignoring pixels with low confidence. We use a subset of 19 labels from Cityscapes.
Our evaluation covers a range of realistic network delays, from 0 to 5 frames, corresponding to 0 to 165 ms. This range is representative of typical round-trip latencies and is sufficient to demonstrate the degradation of conventional remote inference and the resilience of Dedelayed. At training time, the delay $\tau$ was sampled per batch from a uniform distribution over this range. This delay, together with the relevant frames from the video, was fed into the remote model. To replicate real-world usage, we applied compression to the uplink video streams. We chose the resolution and frame rate to fit within reasonable uplink capacity using the lossy WebP image codec at quality 85.
Training details
We adopt a multi-stage training strategy, as detailed in Table˜2. The remote and local models are first trained individually and then later combined. Each model is pretrained on the large-scale ImageNet dataset (Russakovsky et al., 2015) for classification, then on the image segmentation task on Cityscapes (Cordts et al., 2016), before being fine-tuned on the smaller BDD100K driving dataset. We specifically train the remote model to have predictive capability by supplying it with a delay-aware (DA) objective: to predict the labels of future frames, conditioned on a tunable delay. By training in stages, we are able to provide guarantees on baseline remote and local model performance. Finally, the remote and local models are glued together, and the full system is jointly fine-tuned on the delay-aware task of streaming semantic segmentation on video.
We train using cross-entropy loss, the Adan (Xie et al., 2024) optimizer, a trapezoidal cosine learning rate schedule, gradient clipping, and selectively applying discriminative fine-tuning or layer-wise learning rate decay (LLRD) (Howard & Ruder, 2018).
Table 2: Training stages for remote, local, and fusion models.
Model Stage Obj. Data Epochs Res. Freeze
Remote (image only)
SegFormer-B5 pre- DU IN1K, CS
SegFormer-B5 1 DU BDD 15 496
Remote (video-predictive)
EfficientViT-Seg-L1 pre- DU IN1K, CS
EfficientViT-Seg-L1+ViT3D 1 DA BDD 10 496 img-bkbn
EfficientViT-Seg-L1+ViT3D 2 DA BDD 10 496
Local (image only)
MSTransformer2D 1 DU IN1K 320 224
MSTransformer2D 2 DU CS 80 336 bkbn + proj
MSTransformer2D 3 DU CS 80 336 LLRD 0.9
MSTransformer2D 4 DU BDD 15 496 LLRD 0.9
Fusion (local image + remote video-predictive)
MSTransformer2D + EfficientViT-Seg-L1+ViT3D 1 DA BDD 10 480/720 remote img-bkbn + 3D
Stage: pre- = pretrained (external source).
Obj.: DA = delay-aware objective; DU = delay-unaware objective.
Data: IN1K = ImageNet-1K; CS = Cityscapes; BDD = Berkeley DeepDrive 100K.
Freeze: bkbn + proj = backbone & MS-projections frozen; remote img-bkbn + 3D = remote image backbone and remote 3D encoder frozen; LLRD = layer-wise learning rate decay.
6 Results
We compare how various inference systems perform under the effect of communication network latency. Fig.˜6 visualizes the various local-only, remote-only, and fused local+remote methods from different stages of our training. Each baseline is effectively an ablation of our final system.
  • Local image and Remote image inference setups process individual frames in the conventional way, though the remote is susceptible to communication network delay.
  • Remote video has access to past frames of context, but only predicts labels for its present view, and thus fares no better than “remote image”.
  • Remote predictive is fed a tunable delay and sustains accuracy by predicting the future.
  • Local + remote predictive represents a Dedelayed system, and is thus able to further sustain accuracy by merging the remote predictive features with fresh local features.
Figure 6: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames).
Figure 7: Segmentation accuracy (mIoU) versus total latency (milliseconds or frames) for selected local model delays. Points are faded as round-trip latency increases. 4 and 8 ms were interpolated.
While the above results demonstrate the dominance that Dedelayed is capable of, one should also account for the impact of local inference delay. We show this in Fig.˜7 by assessing accuracy versus total latency. For local inference delays of $\leq$8 ms, the “local + remote predictive” method is consistently better across all network round-trip delays in terms of both accuracy and total latency.
7 Conclusion
Dedelayed addresses a central challenge in real-time systems that rely on remote computation: prediction staleness induced by network delay. It mitigates remote inference delay by elevating delay to a first-class variable, conditioning the remote model via a learnable delay embedding, and fusing remote features with fresh local features. Across realistic network conditions, Dedelayed surpasses strong local-only and remote-only baselines, with a particular advantage for longer latencies and high-motion content. As a foundational framework, Dedelayed applies to a wide range of real-time problem domains, enabling intelligent systems that are not only accurate but also truly timely and dependable in dynamic environments. Future work includes studying variable and stochastic delay distributions, high-motion data, lighter local models, and local future prediction.
Appendix A Appendix
A.1 Code
The fused “local + remote predictive” model is defined below:
class FusedModel($n$.Module):
def __init__(self, cls_classes=1000, seg_classes=19):
super().__init__()
self.local_model = MSTransformer2D(cls_classes, seg_classes)
self.remote_model = EfficientViTSeg3D()
self.mlp_pre_pool = PrepoolBlock()
self.mlp_post_pool = PostpoolBlock()
def forward(self, x_local, x_remote, delay):
# Local feature extraction:
h = self.local_model.T1(x_local)
_, _, H, W = h.shape
# Remote feature extraction and pooling:
z = x_remote
z = self.remote_model.forward_features(z, delay)
z = einops.rearrange(z, “b c f h w -> b (c f) h w”, f=4)
z = self.mlp_pre_pool(z)
z = F.adaptive_avg_pool2d(z, output_size=(H, W))
z = self.mlp_post_pool(z)
# Local and remote feature fusion:
h = h + z
# Remaining local model:
y = h
h = self.local_model.T2(h)
y = y + self.local_model.P2(h)
h = self.local_model.T3(h)
y = y + self.local_model.P3(h)
y = self.local_model.seg_head(y)
return y
class EfficientViTSeg3D($n$.Module):
def __init__(self, name=“efficientvit-seg-l1-cityscapes”, seg_classes=19):
super().__init__()
self.image_model = create_efficientvit_seg_model(name)
self.delay_embedding = DelayEmbedding()
self.vit3d = $n$.Sequential(*[
VitBlock3D(in_channels=256, head_dim=32, expand_ratio=4)
for _ in range(12)
self.head = RemotePredictiveHead(seg_classes=seg_classes)
def pool(self, x):
x_pool = F.adaptive_avg_pool3d(x, output_size=(4, x.shape[-2:]))
return einops.rearrange(x_pool, “b c f h w -> b (c f) h w”, f=4)
def forward_features(self, x, delay):
video_embedding = self.image_model.backbone(x) # shape: (B, C, F, H, W)
delay_embedding = self.delay_embedding(delay, video_embedding.shape)
return self.vit3d(video_embedding + delay_embedding)
# For pretraining remote predictive model:
def forward(self, x_remote, delay):
return self.head(self.pool(self.forward_features(x_remote, delay)))
class MSTransformer2D($n$.Module):
“”“Any 2D image segmentation model. We use one with T1, T2, T3 blocks.”“”
To train or evaluate the models, simply feed in frames separated by the appropriate delay and compute cross-entropy loss or mIoU in the typical fashion.
A.2 Effect of delay jitter
We evaluate how our model performs under delay jitter, i.e., when the delay varies for each frame.
Our model is trained only to maximize accuracy for a fixed, tunable delay input and is not explicitly optimized for jitter. Nonetheless, temporal structure in the data allows it to retain accuracy even when the delay input differs from the observed delay. Fig.˜8 characterizes this, showing performance under different observed delays when the model is force-fed a possibly inaccurate delay as input. Accuracy peaks when the delay input matches the observed delay, and the drop is less sharp when the observed delay exceeds the delay input. The latter might be attributed to selection bias—if the predicted features are relevant for a future frame, they often remain useful for subsequent frames.
Figure 8: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Here, the remote model’s delay input has been force-fed a specific value.
In Fig.˜9, we report the above results in matrix form. Although our model was not explicitly trained for mismatched delays or delays beyond 5 frames, it continues to perform well under these conditions. This matrix can also be used to estimate the performance under delay jitter. We model the observed delay as $\tau_{\text{obs}}\sim\mathcal{N}(\mu=\tau_{\text{in}},\,\sigma^{2})$, centered at the delay input to the model, $\tau_{\text{in}}$. The expected accuracy is obtained by taking a weighted sum over mIoU values for each observed delay, using a discretely binned normal probability mass function, with out-of-bounds mass assigned to the boundary bins. The resulting performance is shown in Fig.˜10 for various values of $\sigma$. In many networks, $\sigma=5\text{ ms}$ and $\sigma=15\text{ ms}$ correspond to relatively high jitter. Yet, even for higher values of $\sigma$, our method’s performance maintains its advantage, even assuming optimistically that the competing baselines experience no jitter. This shows that our method performs stably in realistic networks. For comparison, traditional remote inference loses 3.4% mIoU within the first frame of delay, a drop far larger than the degradation our method incurs with the corresponding $\sigma$.
Figure 9: Segmentation accuracy (mIoU) over observed delay and model delay input.
Figure 10: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Latency jitter is modeled as a normal distribution.
A.3 Local input resolution
We evaluate on various local input resolutions. Our remote-assisted local model is able to operate at far lower resolutions without losing accuracy.
Figure 11: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Further fine-tuned and evaluated on various local input resolutions.

### v2

@@
 Abstract
-Remote inference allows lightweight devices to leverage powerful cloud models.
-However, communication network latency makes predictions stale and unsuitable for real-time tasks.
-To address this, we introduce Dedelayed, a delay-corrective method that mitigates arbitrary remote inference delays, allowing the local device to produce low-latency outputs in real time.
-Our method employs a lightweight local model that processes the current frame and fuses in features that a heavyweight remote model computes from past frames.
-On video from the BDD100K driving dataset, Dedelayed improves semantic segmentation accuracy over the stronger of the local-only and remote-only baselines across all realistic communication network delays beyond 33 ms.
-Without incurring additional delay, it improves accuracy by 6.4 mIoU compared to fully local inference and 9.8 mIoU compared to remote inference, for a round-trip delay of 100 ms.
-The advantage grows under longer delays and higher-motion scenes, as delay-mitigated split inference sustains accuracy more effectively, providing clear advantages for real-time tasks that must remain aligned with the current world state.
+Video comprises the vast majority of bits that are generated daily, and is the primary signal driving current innovations in robotics, remote sensing, and wearable technology.
+Yet, the most powerful video understanding models are too expensive for the resource-constrained platforms used in these applications.
+One approach is to offload inference to the cloud; this gives access to GPUs capable of processing high-resolution videos in real time.
+But even with reliable, high-bandwidth communication channels, the combined latency of video encoding, model inference, and round-trip communication prohibits use for certain real-time applications.
+The alternative is to use fully local inference; but this places extreme constraints on computational and power costs, requiring smaller models and lower resolution, leading to degraded accuracy.
+To address these challenges, we propose Dedelayed, a real-time inference system that divides computation between a remote model operating on delayed video frames and a local model with access to the current frame.
+The remote model is trained to make predictions on anticipated future frames, which the local model incorporates into its prediction for the current frame.
+The local and remote models are jointly optimized with an autoencoder that limits the transmission bitrate required by the available downlink communication channel.
+We evaluate Dedelayed on the task of real-time streaming video segmentation using the BDD100k driving dataset.
+For a round trip delay of 100 ms, Dedelayed improves performance by 6.4 mIoU compared to fully local inference and 9.8 mIoU compared to remote inference—an equivalent improvement to using a model ten times larger.
 1 Introduction
-In soft real-time applications—such as cloud gaming or video conferencing—late outputs may be diminished in value, but are still useful.
-In these applications, we can offload expensive computations to powerful cloud GPUs to save on-device power.
+Figure 1: Overview of various inference setups, including conventional local, remote, and split inference.
+Dedelayed combines a small on-device image model and a heavier cloud-based temporally predictive video model to produce accurate and on-time predictions.
+In soft real-time applications—such as cloud gaming or video conferencing [jiang2025real, isikdogan2020eye, jain2025ntire]—late outputs may be diminished in value, but are still useful.
+In these applications, expensive Deep Neural Network (DNN)-based operations can be offloaded to powerful cloud GPUs to save on-device power.
 As long as the typical latency is low, the loss of utility from latency is outweighed by power savings and extended battery life.
-However, in hard real-time applications—such as aerial robotic control or obstacle avoidance—late outputs can be catastrophic, and the system must be designed with a guaranteed deadline.
-Due to the irreducible high-tail latency in wireless communication, hard real-time applications must be equipped with a fully functional local inference pipeline as a fallback in the case that the remote predictions fail to meet the deadline.
-In this work, we focus on such real-time applications running on resource-constrained devices, relying on inference using video inputs.
-In recent years, various approaches to split computing have been proposed to offload computation of expensive image and video models to the cloud to enable the next-generation of robotic, remote sensing, and wearable technology platforms.
+In hard real-time applications—such as aerial robotic control or obstacle avoidance [liu2022large, moon2021real]—late outputs can be catastrophic, and the system must be designed with a guaranteed deadline.
+Due to the irreducible high-tail latency in wireless communication, hard real-time applications must be equipped with a fully functional local inference pipeline as a fallback in the case that the remote predictions fail to meet the deadline [chen2024fogros2, chen2025fogros2].
+Modern hardware for capturing digital video can operate with extremely low power while ingesting super-human amounts of information at high spatial and temporal resolution [engel2023project].
+Using video as the primary or exclusive modality of perception provides an opportunity to significantly reduce size, weight, and power by replacing multiple specialized sensors, and is a major focus of ongoing research.
+Existing video understanding systems based on masked prediction and diffusion-based foundation models [ravi2024sam, agarwal2025cosmos, kim2024openvla] are able to directly produce useful predictions or actions from video input, but require too much computation to run directly on the sensor devices used in these applications.
+Various approaches to split computing and collaborative inference [kang2017neurosurgeon, matsubara2022split, wang2024end, gao2025feature] have been proposed to offload computation of these expensive image and video models to the cloud.
 For real-time streaming video applications, existing approaches still fall into one or more of three common pitfalls.
@@
 (3) They operate on videos with significantly reduced spatiotemporal resolution to manage computational cost, leaving out rich visual details available from modern camera systems.
-Figure 1: Overview of the Dedelayed real-time inference setup. The lightweight local model and a powerful remote model augment each other’s strengths to produce accurate and timely outputs.
-To address these limitations, we introduce Dedelayed (Fig.˜1), a co-inference framework that leverages fresh local information to mitigate the effects of remote inference delay.
-It consists of a local model and a remote model connected over a communication network, and fuses delayed remote signals into a local pipeline operating on fresh on-device inputs, yielding real-time performance that is never worse than either model in isolation.
-This enables the use of high-capacity cloud models for delay-sensitive applications while avoiding the pitfalls of prior offloading approaches:
+To address these limitations, we introduce Dedelayed (Fig.˜1), a collaborative inference framework suitable for hard real-time applications.
+Dedelayed is designed to maximally utilize high-resolution, full-motion video understanding models on cloud GPUs, while avoiding the pitfalls of other split-computing systems:
   1. 1.
@@
 During supervised training of the remote model component, we simulate a delay of $D$ frames.
-In other words, the remote model is trained to predict the future.
+In other words, the remote model is trained to predict features useful in the future.
 A delay embedding—similar to a position embedding in text or vision transformers—allows the behavior of the remote model to adapt to changes in the channel.
@@
     Mixed-resolution inference.
-On-device AI video processing at or near the capture resolution and frame rate—typically >1>1 megapixel and >20>20 frames per second—is rarely feasible, even with lightweight models.
-To save resources, real-time computer vision applications often process each frame independently or use small motion updates instead of natively processing a dense 3D pixel volume.
-Dedelayed is capable of using a mixed-resolution—instead of reducing the resolution for the entire inference pipeline, only the local model resolution is reduced.
-In parallel, the remote model operates on multiple, high-resolution frames using a 3D transformer.
-Thus, the remote model can utilize powerful GPUs to model fine details and motion of delayed frames, while the local component can allocate its resources to modeling the current state of objects and the scene, as shown in Fig.˜3.
+On-device AI video processing at capture resolution and frame rate is rarely feasible, even with lightweight models.
+Dedelayed enables mixed-resolution inference—the local model runs at a lower resolution, and the remote model processes high-resolution frames with a 3D transformer that understands motion.
+Thus, the remote model supplies delayed yet accurate, high-level semantic features, while the local model aligns and localizes them to the current scene, as shown in Fig.˜3.
 Our contributions are threefold:
@@
   2. 2.
-    We introduce Dedelayed, a split computing framework for video inputs that integrates the output of a future-predicting remote model with the current output of a local model.
+    We introduce Dedelayed, a co-inference framework that integrates the output of a future-predicting remote model with the current input to a local model.
   3. 3.
     Using Dedelayed, we create a video segmentation system for urban driving scenes that outperforms any existing local or remote inference solution, while avoiding the pitfalls that limit the practicality of previous approaches.
-For experimental validation, we demonstrate a simple Dedelayed system with an addition-based fusion on off-the-shelf models with minimal architectural changes.
-This makes it easy to enhance existing pipelines, deploy in practice, and extend to other real-time methods.
 Figure 2: To demonstrate the effect of temporally predictive training, we train a 3D transformer to predict the next frame with an MSE loss on pixels.
@@
 2 Background
-In the human visual system, the optic nerve can only transmit a small fraction of the information received by the retina (Kelly, 1962).
-Barlow’s efficient coding hypothesis (Barlow et al., 1961) posits that compression is the primary role of early processing; once this compressed representation is received in deeper layers of the visual cortex, more metabolically intense processing can occur.
-In the predictive coding model (Rao & Ballard, 1999), this processing is driven by feedback mechanisms that minimize a temporally predictive error signal to create a perceptual model that is consistent with sensory inputs.
-Machines equipped with digital video sensors—which are at the center of ongoing innovation in robotics (Kim et al., 2024; O’Neill et al., 2024), remote sensing (Szwarcman et al., 2024; Khani et al., 2021), and wearable technology (Grauman et al., 2022; 2024)—share similar constraints.
-The throughput and power efficiency of ingesting pixels on the sensor device (e.g., a battery-powered robot) are extremely high—typically tens or hundreds of megapixels per watt-second (Engel et al., 2023).
-However, moderately sized DNNs can only process visual data at about one megapixel per watt-second (Cai et al., 2023).
-For more advanced video AI based on autoregressive modeling (Agarwal et al., 2025) or temporal prediction (Assran et al., 2025), the efficiency may be as low as 500 pixels per watt-second.
-Instead of on-device processing, power constraints can be circumvented by compressing and transmitting video streams to cloud GPU datacenters supported by a 100-megawatt power infrastructure (Goldberg & Kehoe, 2013).
-However, fully remote processing is challenging for certain real-time applications (e.g., collision avoidance) due to unreliability in network and cloud infrastructure (Chen et al., 2024; 2025).
-Thus, delivering predictions by a guaranteed deadline requires a fallback procedure independent of the remote server.
-In many systems (e.g., autonomous motor vehicles) the limited accuracy and reliability of lightweight local models warrant a human operator as the fallback (Committee, 2021), preventing full automation.
+Machines equipped with digital video sensors—which are at the center of ongoing innovation in robotics [kim2024openvla, o2024open], remote sensing [szwarcman2024prithvi, khani2021realtimemodelstreaming], and wearable technology [grauman2022ego4d, grauman2024ego]—produce high-bandwidth streams that require significant information processing.
+The throughput and power efficiency of ingesting pixels on the sensor device (e.g., a battery-powered robot) are extremely high—typically tens or hundreds of megapixels per watt-second [engel2023project].
+However, moderately sized DNNs can only process visual data at about one megapixel per watt-second [cai2023efficientvit].
+For more advanced video AI based on autoregressive modeling [agarwal2025cosmos] or temporal prediction [assran2025vjepa2], the efficiency may be as low as 500 pixels per watt-second.
+Instead of on-device processing, power constraints can be circumvented by compressing and transmitting video streams to cloud GPU datacenters supported by a 100-megawatt power infrastructure [goldberg2013cloud].
+Nevertheless, fully remote processing is challenging for certain real-time applications (e.g., collision avoidance) due to unreliability in network and cloud infrastructure [chen2024fogros2, chen2025fogros2].
+This has motivated split computing systems [teerapittayanon2016branchynet, kang2017neurosurgeon, mpegAI2025, choi2018deepfeaturecompression, choi2018nearlosslessdeepfeaturecompression, choi2022scalableimagecoding, azizian2022privacy] which can reduce the bandwidth and latency of remote inference by leveraging both on-device and remote computation.
+Still, delivering predictions by a guaranteed deadline requires a fallback procedure independent of the remote server.
+In many systems (e.g., autonomous motor vehicles) the limited accuracy and reliability of lightweight local models warrant a human operator as the fallback [on2021taxonomy], preventing full automation.
 3 Related work
-Prior work has extensively explored lightweight architectures (e.g., EfficientViT (Cai et al., 2023), MobileNetV4 (Qin et al., 2024)) that minimize computation to achieve real-time on-device performance.
-These approaches deliver low latency but are constrained by device power and compute.
-When devices are too limited, the common alternative is fully remote inference, which offloads computation to servers, but is highly susceptible to network latency.
-Split-computing approaches largely focus on distributing workloads rather than optimizing for strict real-time operation.
-Next-generation compression standards such as MPEG AI (ISO/IEC, 2025) and JPEG AI (Ascenso et al., 2023) target bandwidth reduction via task-specific compression, significantly lowering transmission costs with reasonable compute overhead.
-However, even they lack explicit mechanisms to anticipate or compensate for network delay, leading to stale predictions misaligned with the current world state.
-Other efforts related to our work have also been explored.
-Clockwork Convnets (Shelhamer et al., 2016) reuse stale features to reduce inference latency, but they offer limited temporal reasoning and operate on a single device.
-Accel (Jain et al., 2019) warps heavy-model features forward with optical flow and corrects them with a lightweight model, but is also not intended for across-network operation.
-Adaptive Model Streaming (Khani et al., 2021) streams weight updates from a server to keep a local model fresh, focusing on model adaptation rather than directly mitigating per-frame staleness from communication latency.
-Though less known, Knowledge Boosting (Srinivas et al., 2024) is very closely related to our work.
-Like us, it fuses delayed remote features with a small on-device model, but it assumes a fixed delay.
-We generalize to longer and variable latencies by conditioning on a tunable delay while keeping the design simple and reusable.
+The challenges of real-time video perception have been approached from many angles.
+Previous research can be grouped into four major categories: (1) improving the accuracy-efficiency trade-off of inference, (2) offloading computation from the sensor device to a remote server, (3) temporally predictive video modeling, and (4) dynamically augmenting the on-device model via remote collaboration.
+Improving the accuracy-efficiency trade-off.
+Over the decade since AlexNet [krizhevsky2012alexnet], the trade-off between accuracy and computational efficiency of image-based DNNs has improved significantly.
+This has made local inference (shown in the first column of Fig.˜1) the de facto standard for many computer vision and robotics applications, since the target perception accuracy is often achievable on-device without introducing the additional failure modes or network requirements.
+Still, video models that process dense 3D pixel volumes at high resolution using convolutional or transformer-based DNNs are typically avoided for real-time applications.
+Instead, it is common to apply a lightweight 2D image model to each frame.
+Using recent 2D models like EfficientViT [cai2023efficientvit], high resolution images can be segmented accurately at real-time throughput (greater than 30 fps) using moderately powered hardware (10–100 Watts).
+For more restricted power budgets (less than 10 Watts) the spatial or temporal resolution must be reduced by one or more orders of magnitude to meet real-time deadlines, significantly reducing accuracy.
+Offloading computation.
+Numerous works have explored ways to offload computation of expensive image and video models to the cloud.
+Early exiting [teerapittayanon2016branchynet], split computing [kang2017neurosurgeon], and feature coding for machines (FCM) [mpegAI2025, choi2018deepfeaturecompression, choi2018nearlosslessdeepfeaturecompression, choi2022scalableimagecoding, azizian2022privacy] focus on partitioning DNN layers into two components: one that runs on the sensor device and another that runs in the cloud.
+The overall workflow of these approaches is shown in the third column of Fig.˜1.
+A limitation of these methods and many of their successors is that they allocate local computation towards a prediction pipeline that terminates in the cloud, rather than on the device.
+This may be desirable for applications where the predictions are only used in the cloud (e.g., remote sensing).
+However, if predicted outputs must be available on-device (e.g., autonomous navigation), the delay incurred by round-trip communication leads to stale predictions.
+Even worse, the exclusive allocation of local computational resources to an inference pipeline that necessarily requires a remote server leaves little room for a readily available local fallback mode in the event of late or dropped predicted outputs.
+Temporally predictive video modeling.
+Regardless of DNN size or added network latency, some amount of inference delay is unavoidable.
+Instead of focusing exclusively on reducing the degree of delay, prior works have explored correcting the estimated world state based on the anticipated delay.
+These corrections require some mechanism for temporally predictive video modeling, and can use on-device optical flow estimates [zhang2024neuflow] or readily available motion vectors from an existing video codec [xiao2024task].
+If video is streamed to a sufficiently powerful remote server, generative video world models based on masked prediction [assran2025vjepa2] or diffusion [agarwal2025cosmos] foundation models can provide temporal prediction.
+In each of these cases, existing methods for temporal prediction are repurposed from another application, rather than specialized and optimized for the desired perception task.
+This prevents temporally predictive abilities or behaviors from being learned or optimized with the rest of the DNN-based layers.
+Dynamically augmenting the on-device model via remote collaboration during streaming inference.
+This category of methods improve the on-device model’s inference capabilities by incorporating remote information at runtime (e.g., streaming weights or fusing remote features).
+They are underexplored but important for real-time applications on computationally limited devices.
+Adaptive Model Streaming [khani2021realtimemodelstreaming] streams weight updates to keep the local model tuned to changing scene characteristics.
+Knowledge Boosting [srinivas2024knowledgeboosting] addresses the problem from a different lens—instead of dynamically updating the model weights, it directly fuses in remote features into a small on-device model.
+It requires separate models trained for specific fixed delays, and evaluates on audio data in a controlled local environment with delays up to 48 ms within a local networking context.
+Our method, Dedelayed, is capable of adapting to arbitrary remote inference delays by conditioning the remote model on the measured delay at runtime, enabling dynamic specialization across varying latency.
+Methods within this category are orthogonal to and compatible with many methods from earlier categories, and with each other.
+Our work extends this rarely-explored direction by explicitly inputting dynamic system characteristics such as delay to improve on-device predictions under realistic, variable latency.
 4 Method
-4.1 Delay mitigation framework
 Figure 4: Time progresses left to right.
@@
 The remote model provides accurate outputs, and yet delivers them with delay.
-By careful combination of both subsystems, Dedelayed is able to provide bounded performance guarantees—it is never worse than either local inference or remote inference independently.
+By careful combination of both subsystems, Dedelayed is able to provide bounded performance guarantees.
+When trained appropriately such that the local and remote subsystems deliver the maximum accuracies that they would conventionally—e.g., by freezing the local and remote image backbones during the final joint training of the entire system—it can be guaranteed that Dedelayed is effectively never worse than either local inference or remote inference independently.
 As we will demonstrate later, we are able to glue together the two subsystems in a way that is simple yet effective.
@@
 Dedelayed can be formulated in simple mathematical terms as follows.
-Given a fresh input frame $x_{t}$ at current time $t$, the final prediction $\hat{y}_{t}$ is computed using a lightweight local model, $f_{\text{light}}$, which processes $x_{t}$ along with time-delayed features $z_{t-\tau}$ from a heavyweight remote model, $f_{\text{heavy}}$.
+Given a fresh input frame $x_{t}$ at current time $t$, the final prediction $\hat{y}_{t}$ is computed using a lightweight local model, $f_{\text{local}}$, which processes $x_{t}$ along with time-delayed features $z_{t-\tau}$ from a heavyweight remote model, $f_{\text{remote}}$.
 To produce powerful predictive features, the remote model is conditioned on the delay $\tau$, and processes a short clip of past frames $x_{\leq t-\tau}$ ending at time $t-\tau$.
 This is expressed by the following equations:
-$z_{t-\tau}=f_{\text{heavy}}(\tau,\,x_{\leq t-\tau})$ (1)
-$\hat{y}_{t}=f_{\text{light}}(x_{t},\,z_{t-\tau})$ (2)
-For clarity, the notation is summarized in Table˜1.
+$z_{t-\tau}=f_{\text{remote}}(\tau,\,x_{\leq t-\tau})$ (1)
+$\hat{y}_{t}=f_{\text{local}}(x_{t},\,z_{t-\tau})$ (2)
+For clarity, the notation is summarized in Tab.˜1.
 Fig.˜4 presents a system diagram that demonstrates the fundamental principle we describe, and shows how information propagates through the various subsystems as time progresses.
@@
 $\hat{y}_{t}$ Prediction for time $t$
-$z_{t-\tau}$ Features outputted by heavyweight model run at time $t-\tau$
-$\tau$ Delay in time between the old and current frame at time $t$
-$D$ Delay in frames between the old and current frame at time $t$
-$f_{\text{light}}$ Lightweight model run at time $t$
-$f_{\text{heavy}}$ Heavyweight model run at time $t-\tau$
+$z_{t-\tau}$ Features outputted by remote model run at time $t-\tau$
+$\tau$ Delay in time between the old and current frame
+$D$ Delay in frames between the old and current frame
+$f_{\text{local}}$ Light local model run at time $t$
+$f_{\text{remote}}$ Heavy remote model run at time $t-\tau$
 The entire Dedelayed system is trained end-to-end to minimize a task-specific loss function, $\mathcal{L}_{\text{task}}$, evaluated against the ground truth $y_{t}$ for the current frame.
@@
 In the next section, we detail how we designed a specific implementation to test the Dedelayed framework in action.
-4.2 Design and Implementation
-Dedelayed consists of a lightweight on-device model and a predictive remote model connected over a communication network.
-The remote side computes delay-conditioned features from past inputs and returns them to the device, where they are fused early into the local model running on the current input.
-To provide a guarantee on baseline performance, these components can first be trained to work independently and later fused and trained jointly.
+Figure 5: Dedelayed workflow. The complementary properties of local and remote inference are exploited to produce real-time streaming video understanding that is both accurate and on time.
+5 Design and implementation
+Dedelayed aims to use remote computation to improve the accuracy and robustness of real-time machine vision on resource constrained sensor devices.
+The design of Dedelayed is guided by three key insights.
+(1) Incorporating the outputs of the remote model as side information to the local model eliminates computational redundancy and provides a fallback suitable for hard real-time applications.
+(2) Temporally predictive training can mitigate the impact of latency due to processing and round-trip communication.
+(3) Decoupling inference into a high resolution full-motion video stream and a low-resolution still-image stream exploits the complementary properties of remote and on-device processing.
+Fig.˜5 shows an overview of the design in the context of semantic segmentation of video frames.
 System overview
@@
   2. 2.
-    Each incoming frame is independently encoded into features using a pretrained 2D ViT backbone on the remote. We maintain a context window of the $K$ most recent features.
+    Each incoming frame is fed into a heavyweight remote model $f_{\text{remote}}$.
+To accelerate inference, this model caches and maintains a context window of the $K$ most recent features, which are computed from each incoming frame using a pretrained 2D ViT backbone.
   3. 3.
@@
   5. 5.
-    The lightweight local model runs on a fresh input $x_{t}$, and fuses in the remote features $z_{t-\tau}$.
+    The lightweight local model $f_{\text{local}}$ runs on a fresh input $x_{t}$, and fuses in the remote features $z_{t-\tau}$.
   6. 6.
     The local model finishes decoding the fused representation and outputs labels $\hat{y}_{t}$.
-Remote predictive module.
-Figure Fig.˜5 visualizes the remote component.
-The remote model processes a fixed context of $K$ past frames ending at $t-\tau$.
-Each frame is encoded independently with a 2D ViT—we use EfficientViT-L1 with an effective $8\times 8$ patch.
-The per-frame features are concatenated along the temporal axis and spatially merged into larger $16\times 16$ patches to keep the sequence length comparable.
-(When $K=4$, the sequence length is identical.) A learned delay embedding determined by the delay $\tau$ is added, and the result is then processed by a 3D ViT followed by learned pooling (MLP–pool–MLP) to produce delay-conditioned features $z_{t-\tau}$.
-We pretrain the remote predictive module by attaching a task head to the 3D ViT backbone and training it to predict the target labels $y_{t}$ from inputs up to time $t-\tau$.
-Figure 5: Overview of the remote video predictive model, trained to predict the label at index $n$ from a $K$-frame context ending at nDn-D.
-Local model and fusion.
-The lightweight local model processes the fresh input $x_{t}$.
-We compute first-stage features $h=\texttt{T1}(x_{t})$ and perform early fusion by element-wise addition with the delayed remote features, $h^{\prime}=h+z_{t-\tau}$.
-Both tensors have shape $96\times H/8\times W/8$, so no projection or resizing is required.
-The fused features $h^{\prime}$ are then processed by the remaining local blocks and decoded to the final output $\hat{y}_{t}$.
-If $z_{t-\tau}$ is unavailable, the local model falls back to $h^{\prime}=h$ and proceeds unchanged.
-5 Experiments
-We focus our evaluation of Dedelayed on the task of real-time semantic segmentation of driving scenes, a domain where timely and accurate perception is critical for applications such as autonomous driving and robotics.
-We demonstrate that delay-aware feature fusion can mitigate remote inference latency, sustaining accuracy even when remote predictions are delayed by long communication network latencies.
+Local image model incorporating remote side information
+The local model processes the most recently collected video frame on the sensor device at low resolution, and consists of three learnable modules, shown in the right half of Fig.˜5: (1) pixel-processing 2D convolutional layers (CNN2D), (2) 2D feature-processing convolution and attention layers (CoAt2D), and task specific (e.g., classification or segmentation) MLP layers (L-MLP).
+If present, features from the remote model are added element-wise to the activation map between the CNN2D and the CoAt2D.
+Our local model can operate at a resolution up to 0.34 megapixels ($704\times 480$) while meeting the target latency of 33 ms (a single frame at 30 fps) on our CPU testbed.
+Video compression
+Each frame captured on the sensor device is processed by a lossy image or video codec to allow transmission over a wireless channel.
+In our experiments, we choose the resolution, framerate, and degree of lossy compression to represent a video signal that is transmissible via 5G cellular uplink (30fps 720p content compressed at rates between 1–10 Mbps).
+Latency-aware remote video model
+The remote model operates on high resolution, compressed video frames, and consists of four learnable modules: (1) a 2D vision transformer (ViT2D), (2) a 3D video transformer (ViT3D), (3) task-specific MLP layers (R-MLP), and (4) a dimensionality-reducing autoencoder (DR-AE).
+During the initial training stages, only modules 1–3 (ViT2D, ViT3D, and R-MLP) are used, as shown in Fig.˜6.
+During the final training stage, the R-MLP module is discarded and replaced with the DR-AE to allow joint training of the local and remote components.
+Due to round trip communication and networking, the predictions from the remote model will be delayed by $D$ frames at runtime.
+During training, an artificial delay is applied to the input, but not to the target.
+Thus, the training objective is to predict the target of a future frame.
+During training, we sample a delay uniformly between 0 and 5 frames (up to 167 ms).
+Additionally, we add a learnable “delay embedding” to the input activation maps of the ViT3D—similar to the position or timestep embedding used in diffusion transformers [peebles2023scalable].
+In our experiments, we provide the remote model with four frames of context and aim for the same target latency (33 ms), but on a high power GPU (rather than CPU) testbed.
+Figure 6: Overview of the remote model component as used during pre-training.
+A video sequence $v[n]$ is captured on the sensor device and transmitted to a remote server, incurring a delay of $D$ frames.
+The remote model operates on a fixed context window of $K$ frames.
+A 2D vision transformer (EfficientViT-L1, effective patch size of $8\times 8$ pixels) is applied independently to each of the $K$ input frames.
+The outputs of the 2D transformer are concatenated along the temporal axis, but spatially merged into larger $16\times 16$ patches, thus maintaining a similar sequence length (when $K=4$, the sequence length is identical).
+A learned delay embedding is added, allowing the remaining layers to behave differently based on the expected value of $D$.
+Finally, 3D video transformer layers and task-specific (e.g., segmentation) layers are applied, generating a prediction corresponding to the current frame ($D$ frames in the future after the latest remote input frame).
+Joint prediction with autoencoded resolution adapter.
+After pretraining the remote model for the desired prediction task, its task-specific layers are discarded and replaced with the DR-AE, consisting of adaptive spatial pooling and a channel bottleneck.
+The degree of adaptive spatial pooling is chosen to match the operating resolution of the local model.
+The sample mean is used as the pooling operator.
+The remainder of the DR-AE consists of MLP layers that terminate in a low-resolution, low-channel count activation map suitable for downlink transmission.
+The output of the DR-AE is added element-wise to the output of the pixel-processing CNN2D.
+Then, the local and remote models undergo a final joint training stage.
+6 Evaluation
 Experimental setup
-Our experiments utilize the BDD100K video dataset (Yu et al., 2020), containing video of driving scenes at 30 frames per second (fps).
-Since the dataset does not provide dense segmentation labels for all video frames, we generated pseudo-labels using a pretrained EoMT (Kerssies et al., 2025) model, ignoring pixels with low confidence.
-We use a subset of 19 labels from Cityscapes.
-Our evaluation covers a range of realistic network delays, from 0 to 5 frames, corresponding to 0 to 165 ms.
-This range is representative of typical round-trip latencies and is sufficient to demonstrate the degradation of conventional remote inference and the resilience of Dedelayed.
-At training time, the delay $\tau$ was sampled per batch from a uniform distribution over this range.
-This delay, together with the relevant frames from the video, was fed into the remote model.
-To replicate real-world usage, we applied compression to the uplink video streams.
-We chose the resolution and frame rate to fit within reasonable uplink capacity using the lossy WebP image codec at quality 85.
+We evaluate Dedelayed on the task of real-time semantic segmentation of driving scenes using the BDD100K video dataset [dataset_bdd100k], containing video of driving scenes at 30 frames per second (fps).
+We use the standard 19 label Cityscapes taxonomy [dataset_Cordts2016Cityscapes] for semantic segmentation of urban scenes.
+Since the BDD100K dataset does not provide dense segmentation labels for all video frames, we generate pseudo-labels using two models.
+For the validation set, we use the pre-trained DepthAnything Cityscapes segmentation model [yang2024depth_anything_v1].
+To our knowledge, this is the most accurate publicly available semantic segmentation model for the Cityscapes taxonomy.
+For the training set, we use EoMT [kerssies2025eomtvitsecretlyimagesegmentation], which provides high accuracy, but is significantly faster for labeling 70k training images.
+We evaluate the performance subject to known delays ranging from 0 to 5 frames, corresponding to 0 to 167 ms at 30 fps.
+At training time, the delay $\tau$ is sampled per batch from a uniform distribution over this range.
 Training details
-We adopt a multi-stage training strategy, as detailed in Table˜2.
+We adopt a multi-stage training strategy, as detailed in Tab.˜2.
 The remote and local models are first trained individually and then later combined.
-Each model is pretrained on the large-scale ImageNet dataset (Russakovsky et al., 2015) for classification, then on the image segmentation task on Cityscapes (Cordts et al., 2016), before being fine-tuned on the smaller BDD100K driving dataset.
-We specifically train the remote model to have predictive capability by supplying it with a delay-aware (DA) objective: to predict the labels of future frames, conditioned on a tunable delay.
-By training in stages, we are able to provide guarantees on baseline remote and local model performance.
-Finally, the remote and local models are glued together, and the full system is jointly fine-tuned on the delay-aware task of streaming semantic segmentation on video.
-We train using cross-entropy loss, the Adan (Xie et al., 2024) optimizer, a trapezoidal cosine learning rate schedule, gradient clipping, and selectively applying discriminative fine-tuning or layer-wise learning rate decay (LLRD) (Howard & Ruder, 2018).
-Table 2: Training stages for remote, local, and fusion models.
-Model Stage Obj. Data Epochs Res. Freeze
-Remote (image only)
-SegFormer-B5 pre- DU IN1K, CS
-SegFormer-B5 1 DU BDD 15 496
+Each model is pretrained on the large-scale ImageNet dataset [dataset_ILSVRC15] for classification, then on the image segmentation task on Cityscapes [dataset_Cordts2016Cityscapes], before being fine-tuned on the smaller BDD100K driving dataset.
+We train the remote model to have temporally predictive capability by supplying it with a delay-aware (DA) objective: to predict the labels of future frames conditioned on the degree of delay.
+During the final training stage, the task-specific layers of the MLP are replaced with an entropy bottleneck and resolution adapter that allow its outputs to be added to the intermediate activation map of the local model.
+Using this configuration, the local and remote models are trained jointly for the target scenario of local segmentation incorporating delayed predictions from the remote model.
+The training loss is per-pixel cross-entropy loss.
+We use the Adan [xie2024adan] optimizer, a warmup-stable-decay learning rate schedule, gradient clipping, and selectively applying discriminative fine-tuning or layer-wise learning rate decay (LLRD) [howard2018ulmfit_discriminative_finetuning_LLRD].
+Table 2: Local and remote model components and training setup.
+Stage Local layers Remote layers Data Res.
 Remote (video-predictive)
-EfficientViT-Seg-L1 pre- DU IN1K, CS
-EfficientViT-Seg-L1+ViT3D 1 DA BDD 10 496 img-bkbn
-EfficientViT-Seg-L1+ViT3D 2 DA BDD 10 496
+1 ViT2D IN1K, CS
+2 ViT2D, ViT3D BDD 496
 Local (image only)
-MSTransformer2D 1 DU IN1K 320 224
-MSTransformer2D 2 DU CS 80 336 bkbn + proj
-MSTransformer2D 3 DU CS 80 336 LLRD 0.9
-MSTransformer2D 4 DU BDD 15 496 LLRD 0.9
-Fusion (local image + remote video-predictive)
-MSTransformer2D + EfficientViT-Seg-L1+ViT3D 1 DA BDD 10 480/720 remote img-bkbn + 3D
-Stage: pre- = pretrained (external source).
-Obj.: DA = delay-aware objective; DU = delay-unaware objective.
+3 CNN2D, CoAt2D IN1K 224
+4 CNN2D, CoAt2D CS 336
+5 CNN2D, CoAt2D BDD 496
+Dedelayed
+6 CNN2D, CoAt2D ViT2D, ViT3D BDD 480/720
 Data: IN1K = ImageNet-1K; CS = Cityscapes; BDD = Berkeley DeepDrive 100K.
-Freeze: bkbn + proj = backbone & MS-projections frozen; remote img-bkbn + 3D = remote image backbone and remote 3D encoder frozen; LLRD = layer-wise learning rate decay.
 7 Results
 We compare how various inference systems perform under the effect of communication network latency.
-Fig.˜6 visualizes the various local-only, remote-only, and fused local+remote methods from different stages of our training.
-Each baseline is effectively an ablation of our final system.
+Tab.˜3 shows the segmentation performance (mIoU) for different configurations (local-only, remote-only, and the proposed local+remote system, Dedelayed).
+Each configuration serves as an ablation on the final design.
   • Local image and Remote image inference setups process individual frames in the conventional way, though the remote is susceptible to communication network delay.
@@
   • Remote predictive is fed a tunable delay and sustains accuracy by predicting the future.
-  • Local + remote predictive represents a Dedelayed system, and is thus able to further sustain accuracy by merging the remote predictive features with fresh local features.
+  • Local + remote predictive represents a Dedelayed system, and so is able to further sustain accuracy by merging the remote predictive features with fresh local features.
+As illustrated in Fig.˜7, the performance of conventional remote inference is significantly higher (roughly 5%pt.
+mIoU) than local inference if communication delay is low.
+However, the accuracy degrades rapidly with increasing network delay, reflecting the detriment of using stale predictions for dense prediction tasks involving motion.
+Beyond 2 frames (67 ms) of round-trip latency, remote inference becomes worse than local inference (8%pt.
+mIoU lower at 167 ms).
+This impact is significantly mitigated by adopting a temporal prediction training objective, which extends the utility of remote inference extending to 5 frames (167 ms).
+Finally, the proposed configuration (Dedelayed) incorporates temporally predictive features from the remote model as side information to the local model, and almost completely eliminates the drop-off in accuracy due to delay.
+At the maximum tested delay of 5 frames, Dedelayed provides a 6.7%pt.
+mIoU improvement over the local baseline—a difference that would typically require using a model roughly 10 times larger [cai2023efficientvit, kerssies2025eomtvitsecretlyimagesegmentation].
 Figure 7: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames).
-Figure 7: Segmentation accuracy (mIoU) versus total latency (milliseconds or frames) for selected local model delays. Points are faded as round-trip latency increases. 4 and 8 ms were interpolated.
-While the above results demonstrate the dominance that Dedelayed is capable of, one should also account for the impact of local inference delay.
-We show this in Fig.˜7 by assessing accuracy versus total latency.
-For local inference delays of $\leq$8 ms, the “local + remote predictive” method is consistently better across all network round-trip delays in terms of both accuracy and total latency.
+Tab.˜3 shows the accuracy of various inference systems under various local and remote subsystem delays.
+For local inference delays of $\leq$8 ms, the “local + remote predictive” method consistently delivers better accuracy across all network round-trip delays.
+Table 3: Accuracy (mIoU) given various subsystem delays (ms).
+Local delay Remote delay
+0 ms 33 ms 67 ms 100 ms 133 ms 167 ms
+Remote (video-predictive)
+0.655 0.649 0.640 0.627 0.611 0.596
+Local (image only)
+0 ms 0.588 0.588 0.588 0.588 0.588 0.588
+4 ms 0.585 0.585 0.585 0.585 0.585 0.585
+8 ms 0.582 0.582 0.582 0.582 0.582 0.582
+33 ms 0.562 0.562 0.562 0.562 0.562 0.562
+Dedelayed
+0 ms 0.661 0.660 0.655 0.652 0.646 0.641
+4 ms 0.656 0.655 0.650 0.647 0.642 0.637
+8 ms 0.652 0.650 0.646 0.643 0.637 0.632
+33 ms 0.624 0.619 0.616 0.612 0.607 0.603
+Sub-frame (4 ms, 8 ms) accuracies were lerped between 0 ms and 33 ms.
+Remote jitter
+We also investigate the impact of remote delay jitter on system accuracy.
+LABEL:fig:miou_by_delay_matrix reports accuracy for when the predictive remote model features are conditioned using a different delay than what is later observed.
+Since the local model is stateless, the only thing important to its accuracy is what delay the received remote features were conditioned on.
+In general, the most recently received remote features provide the largest boost to local model accuracy.
+If the remote features are extremely stale, or are missing, the local model may simply ignore them during inference.
 8 Conclusion
@@
 Future work includes studying variable and stochastic delay distributions, high-motion data, lighter local models, and local future prediction.
-Appendix A Appendix
-A.1 Code
-The fused “local + remote predictive” model is defined below:
 class FusedModel($n$.Module):
@@
 class MSTransformer2D($n$.Module):
-“”“Any 2D image segmentation model. We use one with T1, T2, T3 blocks.”“”
-To train or evaluate the models, simply feed in frames separated by the appropriate delay and compute cross-entropy loss or mIoU in the typical fashion.
-A.2 Effect of delay jitter
-We evaluate how our model performs under delay jitter, i.e., when the delay varies for each frame.
-Our model is trained only to maximize accuracy for a fixed, tunable delay input and is not explicitly optimized for jitter.
-Nonetheless, temporal structure in the data allows it to retain accuracy even when the delay input differs from the observed delay.
-Fig.˜8 characterizes this, showing performance under different observed delays when the model is force-fed a possibly inaccurate delay as input.
-Accuracy peaks when the delay input matches the observed delay, and the drop is less sharp when the observed delay exceeds the delay input.
-The latter might be attributed to selection bias—if the predicted features are relevant for a future frame, they often remain useful for subsequent frames.
-Figure 8: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Here, the remote model’s delay input has been force-fed a specific value.
-In Fig.˜9, we report the above results in matrix form.
-Although our model was not explicitly trained for mismatched delays or delays beyond 5 frames, it continues to perform well under these conditions.
-This matrix can also be used to estimate the performance under delay jitter.
-We model the observed delay as $\tau_{\text{obs}}\sim\mathcal{N}(\mu=\tau_{\text{in}},\,\sigma^{2})$, centered at the delay input to the model, $\tau_{\text{in}}$.
-The expected accuracy is obtained by taking a weighted sum over mIoU values for each observed delay, using a discretely binned normal probability mass function, with out-of-bounds mass assigned to the boundary bins.
-The resulting performance is shown in Fig.˜10 for various values of $\sigma$.
-In many networks, $\sigma=5\text{ ms}$ and $\sigma=15\text{ ms}$ correspond to relatively high jitter.
-Yet, even for higher values of $\sigma$, our method’s performance maintains its advantage, even assuming optimistically that the competing baselines experience no jitter.
-This shows that our method performs stably in realistic networks.
-For comparison, traditional remote inference loses 3.4% mIoU within the first frame of delay, a drop far larger than the degradation our method incurs with the corresponding $\sigma$.
-Figure 9: Segmentation accuracy (mIoU) over observed delay and model delay input.
-Figure 10: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Latency jitter is modeled as a normal distribution.
-A.3 Local input resolution
-We evaluate on various local input resolutions. Our remote-assisted local model is able to operate at far lower resolutions without losing accuracy.
-Figure 11: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Further fine-tuned and evaluated on various local input resolutions.
+# Any 2D image segmentation model. We use one with T1, T2, T3 blocks.
+9 Appendix
+9.1 Code
+The fused “local + remote predictive” model is defined above.
+To train or evaluate the models, frames are separated by the appropriate delay and optimized for per-pixel cross-entropy loss or evaluated using mIoU.